Twins Diana and Riana are each 6 years older than their younger brother Kyle. Together, their three ages total 21 years. How old is Kyle?
I know that K+6 =twins
K=Kyle

Answer by rfer(16322)

x+x+6+x+6=21
3x+12=21
3x=9
x=3 Kyle
x+3=9 twins